OpenStudy (anonymous):

[6.03] Solve the following system of equations. x - 2y = 6 2x - 4y = 10 Infinitely Many Solutions No Solutions (0, 0) (6, 10)

jimthompson5910 (jim_thompson5910):

Hint: x - 2y = 6 x = 6+2y -------------- 2x - 4y = 10 2(6+2y) - 4y = 10 12 + 4y - 4y = 10 I'll let you finish

OpenStudy (anonymous):

what is the answer then

jimthompson5910 (jim_thompson5910):

what is 4y-4y

OpenStudy (anonymous):

is 0

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

now what

jimthompson5910 (jim_thompson5910):

so 12 + 4y - 4y = 10 12 + 0y = 10 12 + 0 = 10 12 = 10 is that true?

jimthompson5910 (jim_thompson5910):

is 12 = 10 true?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

12 is equal to 10?

jimthompson5910 (jim_thompson5910):

are you sure?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

12 = 10 is false this means 12 + 4y - 4y = 10 is false for any value of y, so there are no solutions to this equation this means that there are no solutions to the original system of equations

OpenStudy (anonymous):

its is not equal so false

jimthompson5910 (jim_thompson5910):

12 = 10 is false because 12 is not equal to 10 it would be true if you had 12 = 12 or 10 = 10

OpenStudy (anonymous):

so no solutions

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can you help me with this. The graph below is represented by which system of inequalities? -x + 3y < 9 y > -2x + 1 -x + 3y 9 y > -2x + 1 -x + 3y < 9 y -2x + 1 None of these systems represent the graph shown.

jimthompson5910 (jim_thompson5910):

graph each inequality and see if it matches up with what is given here

jimthompson5910 (jim_thompson5910):

I recommend using geogebra to graph the inequalities

OpenStudy (anonymous):

ok